The measure of one angle is three times the measure of its complement. Find the measure of each angle.
Answer by malakumar_kos@yahoo.com(315)


The measure of one angle is three times the measure of its complement. Find the measure of each angle.

Let the required angle be = x
its compliment will be = 90-x
given angle = 3(compliment)
x = 3(90-x)
x = 270-3x
3x+x = 270
4x = 270
x = 270/4 = 67.5 degrees